merchant has coffee worth 50.00 a pound that wishes to mix with 30 pounds of coffee worth 90.00 a pound to get a mixture that can be sold for 60.00 a pound. How many pounds of the 50.00 coffee should be used.
Answer by mananth(16946)

--------price ---------------- quantity
Coffee I 90 ---------------- 30 lb
Coffee II 50 ---------------- x lb
Mixture 60 ---------------- 30 + x lb

90*30+50x=60 ( 30 + x )
2700 + 50 x = 1800 + 60 x
50 x + -60 x = 1800 - -2700
-10 x = -900
/ -10
x = 90
90 lb of 50 % Coffee II